# Ionospheric and Plasmaspheric Delay Characterization for Lunar Terrestrial GNSS Receivers with Global Core Plasma Model

### v1

@@ 1 Introduction
In the past decade, significant progress has been made toward using terrestrial Global Navigation Satellite System (GNSS) signals for positioning, navigation, and timing (PNT) at the Moon. Studies have long shown that GNSS signals can be tracked at lunar distances, despite being far outside their intended coverage. This concept relies on side-lobe signals that spill over beyond Earth, providing a weak but usable navigation source in lunar orbit. Several feasibility analyses [4, 10] used simulations that combine high-sensitivity GNSS receivers and onboard orbit filters, indicating that continuous GNSS navigation from Earth to the Moon is achievable in principle.
@@
Recent on-orbit demonstrations have validated GNSS use near the Moon. In February 2025, NASA and the Italian Space Agency’s Lunar GNSS Receiver Experiment (LuGRE) receiver on a commercial lander successfully acquired GPS and Galileo signals in lunar orbit, which is 401,000 km from the GNSS satellite [36, 24]. During a one-hour tracking period, LuGRE attained a navigation fix, achieving $\sim$1.5 km position accuracy and $\sim$2 m/s velocity accuracy in lunar orbit. In parallel, the European Space Agency is developing the Lunar Pathfinder satellite, slated for launch around 2025–2026, which will be equipped with the NaviMoon receiver that will test GNSS-based navigation in lunar orbit [15].
@@
Two primary challenges limit terrestrial GNSS positioning and timing performance at the Moon: the extreme propagation distance produces very weak signals, and the resulting geometry leads to poor dilution of precision. A range of techniques has been proposed to mitigate these limitations, including time-differenced carrier phase navigation [20], inter-satellite or differential measurements [8, 21, 11], and fusion with complementary sensors [42]. All of these approaches depend critically on minimizing the user equivalent range error (UERE), making accurate modeling of signal delays essential for robust positioning under low-observability conditions.
@@
Among potential error sources, ionospheric and plasmaspheric delays are particularly important. Dual-frequency receivers can cancel most of the ionospheric contribution, but at the cost of increased receiver noise, which can be critical when tracking weak side-lobe signals. Single-frequency terrestrial users typically rely on empirical models such as the Klobuchar model [31] or Galileo’s NTCM-G [19], but these models are not directly applicable to lunar users. Unlike a near-vertical path to the terrestrial users, signals received by lunar users traverse the ionosphere and plasmasphere horizontally, following long paths dominated by plasmaspheric propagation. Accurate delay prediction, therefore, requires explicit modeling of electron density in the plasmasphere.
@@
The plasmasphere, a region of cold, diffuse plasma extending from roughly 1,000 km to beyond 20,000 km altitude, has been the focus of extensive modeling over the past decade. Although its electron density is lower than that of the ionosphere, it can contribute significantly to the total electron content (TEC) along high-altitude or nighttime GNSS paths [30]. Several empirical models are available, including the Global Core Plasma Model (GCPM)[14], IRI-Plas[38], NeQuick [1], and the Neustrelitz Plasmasphere Model (NSPM) [26]. GCPM provides a flexible, density-based description of the plasmasphere; IRI-Plas extends the International Reference Ionosphere to GNSS altitudes; NeQuick offers a time-dependent three-dimensional representation of the topside ionosphere and plasmasphere; and NSPM assimilates ground and satellite data for near–real-time applications.
@@
Previous studies have investigated plasmaspheric delays for GNSS receivers in low-Earth and geosynchronous orbits. For example, [7] developed a global plasmaspheric electron content model using LEO GNSS measurements, and [33] applied first-order delay corrections from GCPM to improve precise orbit determination of a geostationary relay satellite. To the best of our knowledge, however, no prior work has characterized plasmaspheric delays for terrestrial GNSS signals observed in lunar orbit. Understanding how main-lobe and side-lobe signals are affected under different solar and geomagnetic conditions is critical for designing robust lunar PNT systems.
@@
The main contributions of this work are as follows. First, we propose a computationally efficient ray-tracing algorithm that iteratively determines the bent ray path using a Nelder–Mead algorithm [35] without requiring gradient evaluations. Second, we present detailed simulations of first-, second-, and third-order ionospheric and plasmaspheric delays, including bending delays, for GNSS signals received at both the lunar south pole and in lunar orbit. Third, we analyze how these delays vary with signal frequency, geomagnetic activity ($K_{p}$ index), and solar activity (R12 index). This paper expands upon our earlier conference publication [22] by incorporating updated simulations, including the ray-tracing example presented in Section 2.4.
@@ 2.1 Ionosphere and Plasmasphere Simulation with GCPM
The Global Core Plasma Model (GCPM)[14] is an empirical electron–density model of the inner magnetosphere that provides both total electron density and the relative concentrations of major ions. It integrates region-specific submodels for the ionosphere (based on the International Reference Ionosphere, IRI[2]), the plasmasphere, plasmapause, magnetospheric trough, and polar cap, with smooth transitions enforced by hyperbolic-tangent functions. Electron density is specified as a function of the $K_{p}$ geomagnetic index, date (to set solar activity parameters for IRI), and position in solar magnetic (SM) coordinates, defined as follows [33]:
@@ 2.2 Ionospheric and Plasmaspheric Delays
For a right-hand circularly polarized signal, such as GNSS signals, the phase refractive index $n$ and the group refractive index $n_{gr}$ is given by [18]
@@
where $f_{p}=n_{e}e^{2}/(4\pi\epsilon_{0}m)$ and $f_{g}=eB/(2\pi m)$ are the plasma and gyro frequencies, respectively. Above, $e$ is the electron charge, $\epsilon_{0}$ is the permittivity of free space, $m$ is the electron mass, $B$ is the magnetic field strength, $f$ is the frequency of the signal, and $\theta$ is the angle between the ray path and the magnetic field line [16].
@@
where $ds$ is the ray path arc, the quantities $d_{I1}=p/f^{2}$, $d_{I2}=q/f^{3}$, and $d_{I3}=u/f^{4}$ are the first-order, second-order, and third-order group delays, respectively. The integral of the electron density $n_{e}$ along the ray path is called the total electron content (TEC). Since the ray path is curved in the ionosphere and plasmasphere, TEC is split into the line-of-sight (LOS) TEC ($TEC_{LOS}$) and $\Delta TEC_{bend}$, which is the difference between the TEC along the curved ray path and the LOS between the transmitter and receiver. The first-order delay is the dominant term, while the second- and third-order delays are below $1\%$ of the first-order delays, for terrestrial users [17] and low-earth orbit receivers [16].
@@ 2.3 Ray Tracing Algorithm
The propagation of the GNSS signal is governed by the vector eikonal equation [3], as follows:
@@
$\frac{d\bm{\hat{s}}}{ds}$ $=-\frac{\bm{\nabla n}-\bm{\hat{s}}(\bm{\hat{s}}\cdot\bm{\nabla n})}{n}$ (13)
@@
In addition to the differential equations above, the trajectory of the ray is constrained by the boundary conditions at the transmitter and receiver [5].
@@
where $\bm{r}_{tx}(t_{tx})$ is the position of the transmitter at time $t_{tx}$. In order to compute the ray path, we need to solve for the unknown initial ray direction $\bm{\hat{s}_{0}}$ and the ray length $s_{f}$ that satisfies the boundary conditions. Several methods have been proposed to solve this problem, such as the shooting method (homing-in approach) [27, 41] and the variational methods [9].
@@
We first divide the ray path into the plasmasphere section and the vacuum section, based on the observation that the electron density is significantly lower in higher altitude regions above 2 to 3 Earth radii (RE) [14]. We assume that the ray path is straight once it exists beyond this boundary (set as 4 RE from Earth’s center in this work). With this assumption, $s_{f}$ that minimizes the distance between the propagated ray and the receiver can be computed as follows:
@@
    Run the Nelder-Mead optimization algorithm [35] to compute the initial ray direction that minimizes the distance between the propagated final position $\bar{\bm{x}}_{rx}$ and the receiver position $\bm{x}_{rx}$. It has been found that the Nelder-Mead simplex algorithm works reliably when solving for the initial ray direction $\bm{\hat{s}_{0}}$ [17]. Instead of directly optimizing for the direction vector $\bm{\hat{s}_{0}}$, we optimize for the azimuth and elevation angles of the initial ray direction, which turns into a 2-D optimization problem. In the propagation for the optimization, we use the stored $d\bm{\hat{s}}/ds$ at each point to compute the ray path, instead of evaluating the electron density at each point, to avoid calling the GCPM at each iteration.
@@ 3.1.1 Lunar Satellites
We simulated five users in the lunar orbit, which are the LCRNS satellites, with the initial orbital elements as shown in Table 1. The initial states were taken from the reference constellation [37], and the initial epoch was set to March 1st, July 2027. The propagated LCRNS satellite orbits are shown in Figure 6. We also simulated the GPS tracking for a user on the lunar south-pole, resulting in a total of six users in the simulation.
@@ 3.1.2 GNSS Satellites and Tracking Simulations
The GNSS antenna patterns were initialized using the references available online. We used the parameters from the NASA antenna characterization experiment (ACE) study [12] for Block II-F, the original data by Lockheed Martin for Block IIR and IIR-M satellites [32], the United States Coast Guard Navigation Center [13] for Block-III, and Publications Office of the European Union [34] for GALILEO satellites. The antenna patterns of the selected GPS and GALILEO satellites and the GNSS receiver we assumed for the lunar users are shown in Figure 7.
We assumed that the GNSS receiver of the lunar users have a peak gain of 14 dBi, with a half-power beamwidth of 6 degrees. We assumed that the GNSS receiver is able to track the signals from the GPS and GALILEO satellites with a C/N0 of 18 dB-Hz or higher, based on the experimental results of the NaviMoon receiver [40]. The receiver noise errors of the pseudorange measurements are simulated assuming a thermal noise in the Delay Lock Loop (DLL), which depends on the carrier-to-noise ratio of the received signal [29].
@@ 3.1.3 Software
For simulation, we created a simulation software that interfaces the GCPM v2.4 model (implemented in Fortran) with the open-source lunar PNT library, LuPNT [23, 43], which is developed in Python and C++. Several modifications were introduced to the original GCPM code to compute electron density values for recent epochs (from 2015 onward) and accept custom solar activity parameters (R12 index) as inputs. The simulation software computes the GNSS and lunar satellite orbits using LuPNT and performs ray tracing by using the GCPM model to compute electron density values along the ray path. We plan to release the simulation software as an open-source package.
@@ 3.2.1 Results for the Baseline Test Case
The baseline scenario was simulated using electron densities for 1 January 2025, 12:00 UTC, with a simulation span of 45 hours—approximately one orbital period of the LCRNS satellites. This epoch was selected because the GCPM model cannot be run for the 2027/03/01 target date due to unavailable historical parameters. To represent moderate geomagnetic activity, we set the $K_{p}$ index to 3.0 (the actual value at the epoch was 6.667). For solar activity, we used historical measurements of R12 = 167.24 and F10.7 = 210.39, corresponding to high solar activity expected near the peak of Solar Cycle 25. As discussed in Sections 3.2.2 and 3.2.3, the resulting delays are sensitive to both geomagnetic and solar activity levels.
@@ 3.2.3 Effect of Solar Activity (R12 Index)
To assess the influence of solar activity on ionospheric and plasmaspheric delays, we simulated electron densities for five levels of the R12 index: $R12=10,50,100,150,$ and $200$. R12 is the 12-month smoothed sunspot number and serves as a standard indicator of solar activity. In the GCPM model, the IRI submodel requires the F10.7 solar flux and IG12 ionospheric indices, which we derived from R12 using the empirical PHaRLAP relationships [6]:
@@ 3.2.4 Simulation Results for Different Frequencies
where $N_{s}=100$ and $\epsilon_{\rho}\sim\mathcal{N}(0,\sigma_{\rho})$ represents DLL thermal noise. The noise variance $\sigma_{\rho}^{2}$ is derived from the carrier-to-noise ratio ($C/N_{0}$) using the expressions in [28]:

### v2

@@ 1 Introduction
In the past decade, significant progress has been made toward using terrestrial Global Navigation Satellite System (GNSS) signals for positioning, navigation, and timing (PNT) at the Moon. Studies have long shown that GNSS signals can be tracked at lunar distances, despite being far outside their intended coverage. This concept relies on side-lobe signals that spill over beyond Earth, providing a weak but usable navigation source in lunar orbit. Several feasibility analyses [Capuano2017, Delepaut2020] used simulations that combine high-sensitivity GNSS receivers and onboard orbit filters, indicating that continuous GNSS navigation from Earth to the Moon is achievable in principle.
@@
Recent on-orbit demonstrations have validated GNSS use near the Moon. In February 2025, NASA and the Italian Space Agency’s Lunar GNSS Receiver Experiment (LuGRE) receiver on a commercial lander successfully acquired GPS and Galileo signals in lunar orbit, which is 401,000 km from the GNSS satellite [parker2022lugre, insidegnss2025lugre]. During a one-hour tracking period, LuGRE attained a navigation fix, achieving $\sim$1.5 km position accuracy and $\sim$2 m/s velocity accuracy in lunar orbit. In parallel, the European Space Agency is developing the Lunar Pathfinder satellite, slated for launch around 2025–2026, which will be equipped with the NaviMoon receiver that will test GNSS-based navigation in lunar orbit [giordano2022pathfinder].
@@
Two primary challenges limit terrestrial GNSS positioning and timing performance at the Moon: the extreme propagation distance produces very weak signals, and the resulting geometry leads to poor dilution of precision. A range of techniques has been proposed to mitigate these limitations, including time-differenced carrier phase navigation [Iiyama2024TDCPNavigation], inter-satellite or differential measurements [Cheung2024PNT, Iiyama2023Diff, Delepaut2024isl], and fusion with complementary sensors [Vila2025fusion]. All of these approaches depend critically on minimizing the user equivalent range error (UERE), making accurate modeling of signal delays essential for robust positioning under low-observability conditions.
@@
Among potential error sources, ionospheric and plasmaspheric delays are particularly important. Dual-frequency receivers can cancel most of the ionospheric contribution, but at the cost of increased receiver noise, which can be critical when tracking weak side-lobe signals. Single-frequency terrestrial users typically rely on empirical models such as the Klobuchar model [Klobuchar1987] or Galileo’s NTCM-G [Hoque2019], but these models are not directly applicable to lunar users. Unlike a near-vertical path to the terrestrial users, signals received by lunar users traverse the ionosphere and plasmasphere horizontally, following long paths dominated by plasmaspheric propagation. Accurate delay prediction, therefore, requires explicit modeling of electron density in the plasmasphere.
@@
The plasmasphere, a region of cold, diffuse plasma extending from roughly 1,000 km to beyond 20,000 km altitude, has been the focus of extensive modeling over the past decade. Although its electron density is lower than that of the ionosphere, it can contribute significantly to the total electron content (TEC) along high-altitude or nighttime GNSS paths [KLIMENKO2015]. Several empirical models are available, including the Global Core Plasma Model (GCPM)[Gallagher2000GCPM], IRI-Plas[Umut2017IRIPlas], NeQuick [Antonio2012], and the Neustrelitz Plasmasphere Model (NSPM) [Jakowski2018NSPM]. GCPM provides a flexible, density-based description of the plasmasphere; IRI-Plas extends the International Reference Ionosphere to GNSS altitudes; NeQuick offers a time-dependent three-dimensional representation of the topside ionosphere and plasmasphere; and NSPM assimilates ground and satellite data for near–real-time applications.
@@
Previous studies have investigated plasmaspheric delays for GNSS receivers in low-Earth and geosynchronous orbits. For example, [Peng2017] developed a global plasmaspheric electron content model using LEO GNSS measurements, and [Matsumoto2023] applied first-order delay corrections from GCPM to improve precise orbit determination of a geostationary relay satellite. To the best of our knowledge, however, no prior work has characterized plasmaspheric delays for terrestrial GNSS signals observed in lunar orbit. Understanding how main-lobe and side-lobe signals are affected under different solar and geomagnetic conditions is critical for designing robust lunar PNT systems.
@@
The main contributions of this work are as follows. First, we propose a computationally efficient ray-tracing algorithm that iteratively determines the bent ray path using a Nelder–Mead algorithm [Nelder1965Simplex] without requiring gradient evaluations. Second, we present detailed simulations of first-, second-, and third-order ionospheric and plasmaspheric delays, including bending delays, for GNSS signals received at both the lunar south pole and in lunar orbit. Third, we analyze how these delays vary with signal frequency, geomagnetic activity ($K_{p}$ index), and solar activity (R12 index). This paper expands upon our earlier conference publication [iiyama2025plasma] by incorporating updated simulations, including the ray-tracing example presented in Section 2.4.
@@ 2.1 Ionosphere and Plasmasphere Simulation with GCPM
The Global Core Plasma Model (GCPM)[Gallagher2000GCPM] is an empirical electron–density model of the inner magnetosphere that provides both total electron density and the relative concentrations of major ions. It integrates region-specific submodels for the ionosphere (based on the International Reference Ionosphere, IRI[Bilitza2024IRI]), the plasmasphere, plasmapause, magnetospheric trough, and polar cap, with smooth transitions enforced by hyperbolic-tangent functions. Electron density is specified as a function of the $K_{p}$ geomagnetic index, date (to set solar activity parameters for IRI), and position in solar magnetic (SM) coordinates, defined as follows [Matsumoto2023]:
@@ 2.2 Ionospheric and Plasmaspheric Delays
For a right-hand circularly polarized signal, such as GNSS signals, the phase refractive index $n$ and the group refractive index $n_{gr}$ is given by [Jakowski12]
@@
where $f_{p}=n_{e}e^{2}/(4\pi\epsilon_{0}m)$ and $f_{g}=eB/(2\pi m)$ are the plasma and gyro frequencies, respectively. Above, $e$ is the electron charge, $\epsilon_{0}$ is the permittivity of free space, $m$ is the electron mass, $B$ is the magnetic field strength, $f$ is the frequency of the signal, and $\theta$ is the angle between the ray path and the magnetic field line [Hoque2011BendingCorrection].
@@
where $ds$ is the ray path arc, the quantities $d_{I1}=p/f^{2}$, $d_{I2}=q/f^{3}$, and $d_{I3}=u/f^{4}$ are the first-order, second-order, and third-order group delays, respectively. The integral of the electron density $n_{e}$ along the ray path is called the total electron content (TEC). Since the ray path is curved in the ionosphere and plasmasphere, TEC is split into the line-of-sight (LOS) TEC ($TEC_{LOS}$) and $\Delta TEC_{bend}$, which is the difference between the TEC along the curved ray path and the LOS between the transmitter and receiver. The first-order delay is the dominant term, while the second- and third-order delays are below $1\%$ of the first-order delays, for terrestrial users [Hoque2008Estimate] and low-earth orbit receivers [Hoque2011BendingCorrection].
@@ 2.3 Ray Tracing Algorithm
The propagation of the GNSS signal is governed by the vector eikonal equation [Born1980Optics], as follows:
@@
$\frac{d\bm{\hat{s}}}{ds}$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\bm{\nabla n}-\bm{\hat{s}}(\bm{\hat{s}}\cdot\bm{\nabla n})}{n}}$ (13)
@@
In addition to the differential equations above, the trajectory of the ray is constrained by the boundary conditions at the transmitter and receiver [Caruso2023RayTracing].
@@
where $\bm{r}_{tx}(t_{tx})$ is the position of the transmitter at time $t_{tx}$. In order to compute the ray path, we need to solve for the unknown initial ray direction $\bm{\hat{s}_{0}}$ and the ray length $s_{f}$ that satisfies the boundary conditions. Several methods have been proposed to solve this problem, such as the shooting method (homing-in approach) [Jones1975RayTracing, Strangeways2000HomingIn] and the variational methods [Coleman2011RayTracing].
@@
We first divide the ray path into the plasmasphere section and the vacuum section, based on the observation that the electron density is significantly lower in higher altitude regions above 2 to 3 Earth radii (RE) [Gallagher2000GCPM]. We assume that the ray path is straight once it exists beyond this boundary (set as 4 RE from Earth’s center in this work). With this assumption, $s_{f}$ that minimizes the distance between the propagated ray and the receiver can be computed as follows:
@@
    Run the Nelder-Mead optimization algorithm [Nelder1965Simplex] to compute the initial ray direction that minimizes the distance between the propagated final position $\bar{\bm{x}}_{rx}$ and the receiver position $\bm{x}_{rx}$. It has been found that the Nelder-Mead simplex algorithm works reliably when solving for the initial ray direction $\bm{\hat{s}_{0}}$ [Hoque2008Estimate]. Instead of directly optimizing for the direction vector $\bm{\hat{s}_{0}}$, we optimize for the azimuth and elevation angles of the initial ray direction, which turns into a 2-D optimization problem. In the propagation for the optimization, we use the stored $d\bm{\hat{s}}/ds$ at each point to compute the ray path, instead of evaluating the electron density at each point, to avoid calling the GCPM at each iteration.
@@ 3.1.1 Lunar Satellites
We simulated five users in the lunar orbit, which are the LCRNS satellites, with the initial orbital elements as shown in Table 1. The initial states were taken from the reference constellation [LCRNS2025], and the initial epoch was set to March 1st, July 2027. The propagated LCRNS satellite orbits are shown in Figure 6. We also simulated the GPS tracking for a user on the lunar south-pole, resulting in a total of six users in the simulation.
@@ 3.1.2 GNSS Satellites and Tracking Simulations
The GNSS antenna patterns were initialized using the references available online. We used the parameters from the NASA antenna characterization experiment (ACE) study [donaldson2020characterization] for Block II-F, the original data by Lockheed Martin for Block IIR and IIR-M satellites [Marquis2015], the United States Coast Guard Navigation Center [Fischer2022] for Block-III, and Publications Office of the European Union [Menzione2024] for GALILEO satellites. The antenna patterns of the selected GPS and GALILEO satellites and the GNSS receiver we assumed for the lunar users are shown in Figure 7.
We assumed that the GNSS receiver of the lunar users have a peak gain of 14 dBi, with a half-power beamwidth of 6 degrees. We assumed that the GNSS receiver is able to track the signals from the GPS and GALILEO satellites with a C/N0 of 18 dB-Hz or higher, based on the experimental results of the NaviMoon receiver [NavimoonIOD]. The receiver noise errors of the pseudorange measurements are simulated assuming a thermal noise in the Delay Lock Loop (DLL), which depends on the carrier-to-noise ratio of the received signal [kaplan2017understanding].
@@ 3.1.3 Software
For simulation, we created a simulation software that interfaces the GCPM v2.4 model (implemented in Fortran) with the open-source lunar PNT library, LuPNT [Iiyama_LuPNT, Vila2025LuPNT], which is developed in Python and C++. Several modifications were introduced to the original GCPM code to compute electron density values for recent epochs (from 2015 onward) and accept custom solar activity parameters (R12 index) as inputs. The simulation software computes the GNSS and lunar satellite orbits using LuPNT and performs ray tracing by using the GCPM model to compute electron density values along the ray path. We plan to release the simulation software as an open-source package.
@@ 3.2.1 Results for the Baseline Test Case
The baseline scenario was simulated using electron densities for 1 January 2025, 12:00 UTC, with a simulation span of 45 hours—approximately one and a half orbital periods of the LCRNS satellites. This epoch was selected because the GCPM model cannot be run for the 2027/03/01 target date due to unavailable historical parameters. To represent moderate geomagnetic activity, we set the $K_{p}$ index to 3.0 (the actual value at the epoch was 6.667). For solar activity, we used historical measurements of R12 = 167.24 and F10.7 = 210.39, corresponding to high solar activity expected near the peak of Solar Cycle 25. As discussed in Sections 3.2.2 and 3.2.3, the resulting delays are sensitive to both geomagnetic and solar activity levels.
@@ 3.2.3 Effect of Solar Activity (R12 Index)
To assess the influence of solar activity on ionospheric and plasmaspheric delays, we simulated electron densities for five levels of the R12 index: $R12=10,50,100,150,$ and $200$. R12 is the 12-month smoothed sunspot number and serves as a standard indicator of solar activity. In the GCPM model, the IRI submodel requires the F10.7 solar flux and IG12 ionospheric indices, which we derived from R12 using the empirical PHaRLAP relationships [pharlap]:
@@ 3.2.4 Simulation Results for Different Frequencies
where $N_{s}=100$ and $\epsilon_{\rho}\sim\mathcal{N}(0,\sigma_{\rho})$ represents DLL thermal noise. The noise variance $\sigma_{\rho}^{2}$ is derived from the carrier-to-noise ratio ($C/N_{0}$) using the expressions in [kaplan2005understanding]: